Identify the maximum number of relative extrema g(x)=2/x^2+3
1
Expert's answer
2019-07-26T09:46:31-0400

In order to find the extremum of a function, you need to find its derivative.


"g(x)=\frac{2}{x^2}+3;"

"\frac{dg(x)}{dx}=2*x^{(-2-1)}*(-2)+0;""\frac{dg(x)}{dx}=\frac{-4}{x^3} \neq 0"

As can be seen, the derivative of the function is not zero for any x

Answer: The function has no extrema.